OpenStudy (anonymous):

find the 100th number in the list 7,12,17

OpenStudy (anonymous):

Find the difference. \[17-12=5, 12-7=5, d=5\] Use the equation: \[t_n=a+(n-1)d\] \[t_100=7+(100-1)5\]